# Stochastic SIR model with individual heterogeneity and infection-age dependent infectivity on large non-homogeneous random graphs

### v1

@@ 3.1. Discussions of the main result
For example, we may consider the $(X_{i}^{N})$ as spatial coordinates, still uniformly sampled in $[0,1]$ yet with the norm-distance $d(x,y)=|x-y|$ on the circle so that all the locations are equivalent (i.e., on the one-dimensional torus with 00 and $1$ identified). Let us then distinguish in the expressions of $\kappa^{N}$ and $\gamma^{N}$ a local interaction pattern (specified by the index $\mathcal{L}$) with radius $\delta\in(0,1/2)$ and a global interaction pattern (with index $\mathcal{G}$) over the whole domain:
@@
[…]
@@ 4.1. Existence and uniqueness of solution to the limiting equations
Under Assumptions 2.2 and 3.3, the set of equations (3.6)–(3.7) has a unique solution $(\bar{\mu}^{S}_{\cdot},\overline{\mathfrak{F}}_{\cdot})$.
@@
We first prove the uniqueness. Before proceeding, we establish some useful bounds. Suppose that $(\bar{\mu}^{S}_{\cdot},\overline{\mathfrak{F}}_{\cdot})$ is a solution. By Assumptions 2.2 and 3.3, we have $\bar{\lambda}(x,t)\leq\lambda^{*}$ and $\bar{\omega}(x,x^{\prime})\leq\omega^{*}$ for some $\omega^{*}>0$. Recalling (3.7) and Assumption 3.1, we derive
$\begin{split}\overline{\mathfrak{F}}(t,x)&\leq\lambda^{*}\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\int_{0}^{\infty}\bigg{[}1+\frac{F_{x^{\prime}}(a^{\prime})}{F^{c}_{x^{\prime}}(a^{\prime})}\bigg{]}\bar{\mu}_{0}^{I}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime})\\ &\quad+\lambda^{*}\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\int_{0}^{t}\overline{\mathfrak{F}}(s,x^{\prime})\exp\bigg{(}-\int_{0}^{s}\overline{\mathfrak{F}}(r,x^{\prime})\mathrm{d}r\bigg{)}\mathrm{d}s\;\bar{\mu}^{S}_{0}(\mathrm{d}x^{\prime})\,\\ &\leq\lambda^{*}\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\cdot\Big{[}\langle\bar{\mu}_{0}^{I}(\mathrm{d}x^{\prime},.),\mathbbm{1}\rangle+\bar{\mu}_{0}^{R}(\mathrm{d}x^{\prime})+\bar{\mu}^{S}_{0}(\mathrm{d}x^{\prime})\Big{]}\\ &\leq\lambda^{*}\omega^{*}<\infty,\end{split}$ (4.1)
@@
Suppose now that there are two solutions $(\bar{\mu}^{S,\ell}_{\cdot},\overline{\mathfrak{F}}^{\ell}_{\cdot})$, $\ell=1,2$. By (3.6) and (3.7), for $\varphi\in C_{b}({\mathbb{R}}_{+})$, we obtain
$\langle\bar{\mu}^{S,1}_{t}-\bar{\mu}^{S,2}_{t},\varphi\rangle=-\int_{0}^{t}\langle\bar{\mu}^{S,1}_{s}-\bar{\mu}^{S,2}_{s},\overline{\mathfrak{F}}^{1}(s)\varphi\rangle\mathrm{d}s-\int_{0}^{t}\langle\bar{\mu}^{S,2}_{s},(\overline{\mathfrak{F}}^{1}(s)-\overline{\mathfrak{F}}^{2}(s))\varphi\rangle\mathrm{d}s\,,$ (4.3)
@@
$\begin{split}\overline{\mathfrak{F}}^{1}(t,x)-\overline{\mathfrak{F}}^{2}(t,x)&=\int_{0}^{t}\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\bar{\lambda}(x^{\prime},t-s)\overline{\mathfrak{F}}^{1}(s,x^{\prime})(\bar{\mu}^{S,1}_{s}(\mathrm{d}x^{\prime})-\bar{\mu}^{S,2}_{s}(\mathrm{d}x^{\prime}))\mathrm{d}s\\ &\qquad+\int_{0}^{t}\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\bar{\lambda}(x^{\prime},t-s)(\overline{\mathfrak{F}}^{1}(s,x^{\prime})-\overline{\mathfrak{F}}^{2}(s,x^{\prime}))\bar{\mu}^{S,2}_{s}(\mathrm{d}x^{\prime})\mathrm{d}s\,.\end{split}$ (4.4)
Let
$D_{t}=(\lambda^{*}\omega^{*})\cdot\left\|\bar{\mu}^{S,1}_{t}-\bar{\mu}^{S,2}_{t}\right\|_{TV}+\|\overline{\mathfrak{F}}^{1}(t,.)-\overline{\mathfrak{F}}^{2}(t,.)\|_{\infty}.$
By (4.3), (4.4), (4.1) and (4.2):
$D_{t}\leq C\int_{0}^{t}D_{s}\mathrm{d}s,$ (4.5)
where $C=2\lambda^{*}\omega^{*}$. Thanks to Gronwall’s inequality, the proof of the uniqueness is concluded.
Finally, the existence can be proved by Picard iteration (which will invoke similar estimates as above). ∎
@@ 4.2. Regularity of $\overline{\mathfrak{F}}(\cdot)$
The function $:x\mapsto\overline{\mathfrak{F}}(.,x)$ with values in the set of bounded measurable functions from $\mathbb{R}_{+}$ to itself equipped with the uniform norm topology is $\bar{\mu}_{X}$-almost everywhere continuous.
@@
Recalling (3.7), with the uniform upper-bound of $\bar{\lambda}$ by $\lambda^{*}$, we first deduce the following inequality for any $t\geq 0$ and any $x_{1},x_{2}\in\mathbf{\mathbb{X}}$:
$\Big{|}\overline{\mathfrak{F}}(t,x_{1})-\overline{\mathfrak{F}}(t,x_{2})\Big{|}\leq\lambda^{*}\int_{\mathbf{\mathbb{X}}}|\bar{\omega}(x_{1},x^{\prime})-\bar{\omega}(x_{2},x^{\prime})|\Big{\{}\Big{(}1+\frac{F_{x^{\prime}}(a^{\prime})}{F^{c}_{x^{\prime}}(a^{\prime})}\Big{)}\bar{\mu}_{0}^{I}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime})\\ +\int_{0}^{t}\overline{\mathfrak{F}}(s,x^{\prime})\exp\Big{(}-\int_{0}^{s}\overline{\mathfrak{F}}(r,x^{\prime})\mathrm{d}r\Big{)}\mathrm{d}s\;\bar{\mu}_{0}^{S}(\mathrm{d}x^{\prime})\Big{\}}.$
@@
As stated in Appendix A in Proposition A.2 and proved just afterwards, the function $x\mapsto\bar{\omega}(x,.)$ is $\bar{\mu}_{X}$ a.e. continuous with the $L^{1}(\bar{\mu}_{X})$ distance, provided that $\bar{\omega}$ is $\bar{\mu}_{X}^{\otimes 2}$-a.e. continuous as in Assumption 3.3. This concludes the proof of Lemma 4.2. ∎
@@ 4.3. Alternative representation of the limit as PDEs
Assume that $F_{x}$ is absolutely continuous with density $f_{x}$ for each $x\in\mathbf{\mathbb{X}}$ and that $F^{c}_{x}(a)>0$ for any $a\in\mathbb{R}_{+}$. Assume that $h_{x}(a)=f_{x}(a)/F^{c}_{x}(a)$, the hazard rate function, is continuous and bounded uniformly in both $x\in\mathbf{\mathbb{X}}$ and $a\in\mathbb{R}_{+}$. Assume moreover that $\bar{\mu}^{I}_{0}(\mathbf{\mathbb{X}}\times\{0\})=0$. Then, $\bar{\mu}^{I}_{t}(\mathrm{d}x,\mathrm{d}a)$ in (3.8) is the unique solution to the following equation, defined for any $\psi\in C^{0,1}(\mathbf{\mathbb{X}}\times{\mathbb{R}}_{+})$ and $t>0$,
@@
$\langle\partial_{t}\bar{\mu}^{I}_{t},\psi\rangle+\langle\partial_{a}\bar{\mu}^{I}_{t},\psi\rangle=-\langle\bar{\mu}^{I}_{t},h\psi\rangle$ (4.8)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle\bar{\mu}^{I}_{t},\psi\rangle$ $=\int_{0}^{\infty}\int_{\mathbf{\mathbb{X}}}\partial_{a}\psi(x,a+t)\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\bar{\mu}_{0}^{I}(\mathrm{d}x,\mathrm{d}a)$
$\quad-\int_{0}^{\infty}\int_{\mathbf{\mathbb{X}}}\psi(x,a+t)\frac{f_{x}(a+t)}{F^{c}_{x}(a)}\bar{\mu}_{0}^{I}(\mathrm{d}x,\mathrm{d}a)$
@@
$\quad-\int_{0}^{t}\int_{\mathbf{\mathbb{X}}}\psi(x,t-s)f_{x}(t-s)\overline{\mathfrak{F}}(s,x)\bar{\mu}^{S}_{s}(\mathrm{d}x)\mathrm{d}s$
$\quad+\int_{0}^{t}\int_{\mathbf{\mathbb{X}}}\partial_{a}\psi(x,t-s)F^{c}_{x}(t-s)\overline{\mathfrak{F}}(s,x)\bar{\mu}^{S}_{s}(\mathrm{d}x)\mathrm{d}s\,.$
@@
$\langle\bar{\mu}^{I}_{t},\partial_{a}\psi\rangle$ $=\int_{0}^{\infty}\int_{\mathbf{\mathbb{X}}}\partial_{a}\psi(x,a+t)\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\bar{\mu}_{0}^{I}(\mathrm{d}x,\mathrm{d}a)$
$\quad+\int_{0}^{t}\int_{\mathbf{\mathbb{X}}}\partial_{a}\psi(x,t-s)F^{c}_{x}(t-s)\overline{\mathfrak{F}}(s,x)\bar{\mu}^{S}_{s}(\mathrm{d}x)\mathrm{d}s\,,$
@@
$\langle\bar{\mu}^{I}_{t},h\psi\rangle$ $=\int_{0}^{\infty}\int_{\mathbf{\mathbb{X}}}\psi(x,a+t)\frac{f_{x}(a+t)}{F^{c}_{x}(a)}\bar{\mu}_{0}^{I}(\mathrm{d}x,\mathrm{d}a)$
$\quad+\int_{0}^{t}\int_{\mathbf{\mathbb{X}}}\psi(x,t-s)f_{x}(t-s)\overline{\mathfrak{F}}(s,x)\bar{\mu}^{S}_{s}(\mathrm{d}x)\mathrm{d}s\,.$
@@
[…]

### v2

@@ 3. Functional law of large numbers
For example, we may consider the $(X_{i}^{N})$ as spatial coordinates, still uniformly distributed in $[0,1]$ yet with the norm-distance d(x,y)=|xy|d(x,y)=|x-y| on the circle so that all the locations are equivalent (i.e., on the one-dimensional torus with 0 and $1$ identified). Let us then distinguish in the expressions of $\kappa^{N}$ and $\gamma^{N}$ a local interaction pattern (specified by the index $\mathcal{L}$) with radius $\delta\in(0,1/2)$ and a global interaction pattern (with index $\mathcal{G}$) over the whole domain:
@@
Then, $\Upsilon^{N}\sim\bar{\omega} \cdot(\bar{\sigma}^{N})^{2}/(3\gamma^{N})$ converges to zero as required in Assumption 3.4 if and only if $(\bar{\sigma}^{N})^{2}\ll\gamma_{N}$. Since $\bar{\sigma}^{N}<\gamma_{N}$ in this case (to keep $\omega^{N}$ non-negative), both properties hold true if and only if $\gamma_{N}$ tends to zero, which is exactly the first condition in Assumption 3.4. So we do not have any additional restriction on $\bar{\sigma}_{N}$ in this model.
@@ 4.1. Existence and uniqueness of solution to the limiting equations
Under Assumptions 2.2 and 3.2, the set of equations (3.5)–(3.6) has a unique solution $(\bar{\mu}^{S}_{\cdot},\overline{\mathfrak{F}}_{\cdot})$.
@@
We first prove the uniqueness. Before proceeding, we establish some useful bounds. Suppose that $(\bar{\mu}^{S}_{\cdot},\overline{\mathfrak{F}}_{\cdot})$ is a solution. By Assumptions 2.2 and 3.2, we have $\bar{\lambda}(x,t)\leq\lambda^{*}$ and $\bar{\omega}(x,x^{\prime})\leq\omega^{*}$ for some $\omega^{*}>0$. Recalling (3.6) and Assumption 2.6, we derive
$\begin{split}\overline{\mathfrak{F}}(t,x)&\leq\lambda^{*}\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\int_{0}^{\infty}\bigg[\frac{F^{c}_{x^{\prime}}(a^{\prime}+t)}{F^{c}_{x^{\prime}}(a^{\prime})}\bigg]\bar{\mu}_{0}^{I}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime})\\ &\quad+\lambda^{*}\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\int_{0}^{t}\overline{\mathfrak{F}}(s,x^{\prime})\exp\bigg(-\int_{0}^{s}\overline{\mathfrak{F}}(r,x^{\prime})\mathrm{d}r\bigg)\mathrm{d}s\;\bar{\mu}^{S}_{0}(\mathrm{d}x^{\prime})\,\\ &\leq\lambda^{*}\omega^{*}\int_{\mathbf{\mathbb{X}}}\Big[\langle\bar{\mu}_{0}^{I}(\mathrm{d}x^{\prime},.),\mathbbm{1}\rangle+\bar{\mu}^{S}_{0}(\mathrm{d}x^{\prime})\Big]\\ &\leq\lambda^{*}\omega^{*}<\infty,\end{split}$ (4.1)
@@
Suppose now that there are two solutions $(\bar{\mu}^{S,\ell}_{\cdot},\overline{\mathfrak{F}}^{\ell}_{\cdot})$, $\ell=1,2$. Let $D_{t}=\|\overline{\mathfrak{F}}^{1}(t,.)-\overline{\mathfrak{F}}^{2}(t,.)\|_{\infty}$. By (4.5):
$\begin{split}\overline{\mathfrak{F}}^{1}(t,x)-\overline{\mathfrak{F}}^{2}(t,x)&=\int_{0}^{t}\hskip-5.69046pt\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\bar{\lambda}(x^{\prime},t-s)\overline{\mathfrak{F}}^{1}(s,x^{\prime})(\bar{\mu}^{S,1}_{s}(\mathrm{d}x^{\prime})-\bar{\mu}^{S,2}_{s}(\mathrm{d}x^{\prime}))\mathrm{d}s\\ &\qquad+\int_{0}^{t}\hskip-5.69046pt\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\bar{\lambda}(x^{\prime},t-s)(\overline{\mathfrak{F}}^{1}(s,x^{\prime})-\overline{\mathfrak{F}}^{2}(s,x^{\prime}))\bar{\mu}^{S,2}_{s}(\mathrm{d}x^{\prime})\mathrm{d}s\,.\end{split}$ (4.3)
Since $a\in\mathbb{R}_{+}\mapsto e^{-a}$ is 1-Lischitz continuous:
$\Big\|\exp\bigg(-\int_{0}^{s}\overline{\mathfrak{F}}^{1}(r,\,.)\mathrm{d}s\bigg)-\exp\bigg(-\int_{0}^{s}\overline{\mathfrak{F}}^{2}(r,\,.)\mathrm{d}r\bigg)\Big\|_{\infty}\leq\int_{0}^{s}D_{r}\,\mathrm{d}r\,.$
We then deduce from (3.9) that: $\left\|\bar{\mu}^{S,1}_{s}-\bar{\mu}^{S,2}_{s}\right\|_{TV}\leq\int_{0}^{s}D_{r}\,\mathrm{d}r$. By injecting this inequality in (4.3) with (4.1) and (4.2), we deduce for any $T>0$ and $t\in[0,T]$ that $D_{t}\leq C\int_{0}^{t}D_{s}\mathrm{d}s$ where $C=\lambda^{*}\omega^{*}+(\lambda^{*}\omega^{*})^{2}\,T$. Thanks to Gronwall’s inequality (for any $T>0$), the proof of the uniqueness is concluded.
Finally, existence can be proved by the following modified Picard iteration. We initialize with $\bar{\mu}^{S,0}_{s}(\mathrm{d}x^{\prime})=\bar{\mu}^{S}_{0}(\mathrm{d}x^{\prime})$, $\overline{\mathfrak{F}}^{0}(t,x)=0$, then for any $k\geq 0$:
$\bar{\mu}^{S,k+1}_{t}(\mathrm{d}x)=\bar{\mu}^{S}_{0}(\mathrm{d}x)-\int_{0}^{t}\overline{\mathfrak{F}}^{k}(s,x)\bar{\mu}^{S,k+1}_{s}(\mathrm{d}x)\mathrm{d}s\,$, (4.4)
@@
$\begin{split}\overline{\mathfrak{F}}^{k+1}(t,x)&=\int_{0}^{\infty}\hskip-8.5359pt\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\frac{\bar{\lambda}(x^{\prime},a^{\prime}+t)}{F^{c}_{x^{\prime}}(a^{\prime})}\bar{\mu}_{0}^{I}(\mathrm{d}x^{\prime},da^{\prime})\\ &\quad+\int_{0}^{t}\hskip-5.69046pt\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\bar{\lambda}(x^{\prime},t-s)\overline{\mathfrak{F}}^{k}(s,x^{\prime})\bar{\mu}^{S,k+1}_{s}(\mathrm{d}x^{\prime})\mathrm{d}s\,.\end{split}$ (4.5)
In these two definitions, we make the choice of the product between $\overline{\mathfrak{F}}^{k}$ and $\bar{\mu}^{S,k+1}$ because then we have the following explicit expression of $\bar{\mu}^{S,k+1}$ in terms of $\overline{\mathfrak{F}}^{k}$:
$\bar{\mu}^{S,k+1}_{t}(\mathrm{d}x)=\exp\bigg(-\int_{0}^{t}\overline{\mathfrak{F}}^{k}(r,x)\mathrm{d}r\bigg)\bar{\mu}^{S}_{0}(\mathrm{d}x)\,$, (4.6)
similarly to (3.9). We can thus justify similarly the a priori estimates (4.2) and (4.1) for the above approximating sequence. Then, the arguments used for proving uniqueness yield the convergence of our Picard iteration, exactly as in the usual case of globally Lipschitz coefficients. ∎
@@ 4.2. Regularity of $\overline{\mathfrak{F}}(\cdot)$
The function $x\mapsto\overline{\mathfrak{F}}(.,x)$ with values in the set of bounded measurable functions from $\mathbb{R}_{+}$ to itself equipped with the uniform norm topology is $\bar{\mu}_{X}$-a.e. continuous.
@@
Recalling (3.6) and Assumption 2.6, we first deduce the following inequality for any $t\geq 0$ and any $x_{1},x_{2}\in\mathbf{\mathbb{X}}$:
$\Big|\overline{\mathfrak{F}}(t,x_{1})-\overline{\mathfrak{F}}(t,x_{2})\Big|\leq\lambda^{*}\int_{\mathbf{\mathbb{X}}}|\bar{\omega}(x_{1},x^{\prime})-\bar{\omega}(x_{2},x^{\prime})|\Big\{\Big(\frac{F^{c}_{x^{\prime}}(a^{\prime}+t)}{F^{c}_{x^{\prime}}(a^{\prime})}\Big)\bar{\mu}_{0}^{I}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime})\\ +\int_{0}^{t}\overline{\mathfrak{F}}(s,x^{\prime})\exp\Big(-\int_{0}^{s}\overline{\mathfrak{F}}(r,x^{\prime})\mathrm{d}r\Big)\mathrm{d}s\;\bar{\mu}_{0}^{S}(\mathrm{d}x^{\prime})\Big\}.$
@@
As stated in Appendix A in Lemma A.2 and proved just afterwards, the function $x\mapsto\bar{\omega}(x,.)$ is $\bar{\mu}_{X}$ a.e. continuous with the $L^{1}(\bar{\mu}_{X})$ distance, since $\bar{\omega}$ is $\bar{\mu}_{X}^{\otimes 2}$-a.e. continuous by virtue of Assumption 3.2. This concludes the proof of Lemma 4.2. ∎
@@ 4.3. Alternative representation of the limit as PDEs
Assume that $F_{x}$ is absolutely continuous with density $f_{x}$ for each $x\in\mathbf{\mathbb{X}}$ and that $F^{c}_{x}(a)>0$ for any $a\in\mathbb{R}_{+}$. Assume that $h_{x}(a)=f_{x}(a)/F^{c}_{x}(a)$, the hazard rate function, is continuous and bounded uniformly in both $x\in\mathbf{\mathbb{X}}$ and $a\in\mathbb{R}_{+}$. Assume moreover that $\bar{\mu}^{I}_{0}(\mathbf{\mathbb{X}}\times\{0\})=0$. Then, $\bar{\mu}^{I}_{t}(\mathrm{d}x,\mathrm{d}a)$ in (3.7) is the unique solution to the following equation: for any $\psi\in C^{0,1}(\mathbf{\mathbb{X}}\times{\mathbb{R}}_{+})$ and $t>0$,
@@
$\langle\partial_{t}\bar{\mu}^{I}_{t},\psi\rangle+\langle\partial_{a}\bar{\mu}^{I}_{t},\psi\rangle=-\langle h\,\bar{\mu}^{I}_{t},\psi\rangle$ (4.9)
@@
We directly identify from Proposition 4.3 and (3.10) that the system of equations under consideration is a weak formulation of the following system. Here, $u^{I}_{t}(x,a)$ (resp. $u^{S}_{t}(x)$) represent the density of $\bar{\mu}_{t}^{I}$ (resp. $\bar{\mu}_{t}^{S}$) with respect to $\bar{\mu}_{X}(\mathrm{d}x)\,\mathrm{d}a$ (resp. $\bar{\mu}_{X}(\mathrm{d}x)$):
$\left\{\begin{split}&\partial_{t}u^{I}_{t}(x,a)+\partial_{a}u^{I}_{t}(x,a)=-h(x,a)\,u^{I}_{t}(x,a)\,,\\ &u^{I}_{t}(x,0)=-\partial_{t}u^{S}_{t}(x)=\overline{\mathfrak{F}}(t,x)\,u^{S}_{t}(x)\,,\\ &\overline{\mathfrak{F}}(t,x)=\int_{0}^{\infty}\hskip-8.5359pt\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\frac{\bar{\lambda}(x^{\prime},a^{\prime})}{F^{c}_{x^{\prime}}(a^{\prime})}u^{I}_{t}(x^{\prime},a^{\prime})\,\bar{\mu}_{X}(\mathrm{d}x^{\prime})\,\mathrm{d}a^{\prime}\,.\end{split}\right$. (4.10)
We can associate to this system the density process $u^{R}$ of recovered individual, defined by $u^{R}_{t}(x)=1-u^{S}_{t}(x)-\int_{0}^{\infty}u^{I}_{t}(x,a)\,\mathrm{d}a$ and solution to $\partial_{t}u^{R}_{t}(x)=\int_{0}^{\infty}h(x,a)\,u^{I}_{t}(x,a)\mathrm{d}a$.
While the uniqueness of the system in Proposition 4.3 entails the one of the above system (by defining $\bar{\mu}^{S}_{t}(\mathrm{d}x)=u^{I}_{t}(x,a)\,\bar{\mu}_{X}(\mathrm{d}x)$ and $\bar{\mu}^{I}_{t}(\mathrm{d}x,\mathrm{d}a)=u^{I}_{t}(x,a)\,\bar{\mu}_{X}(\mathrm{d}x)\,\mathrm{d}a$), the regularity of such solutions and in what stronger sense the system (4.10) is satisfied is left open.
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle\bar{\mu}^{I}_{t},\psi\rangle$ $=\int_{0}^{\infty}\hskip-8.5359pt\int_{\mathbf{\mathbb{X}}}\partial_{a}\psi(x,a+t)\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\bar{\mu}_{0}^{I}(\mathrm{d}x,\mathrm{d}a)$
$\quad-\int_{0}^{\infty}\hskip-8.5359pt\int_{\mathbf{\mathbb{X}}}\psi(x,a+t)\frac{f_{x}(a+t)}{F^{c}_{x}(a)}\bar{\mu}_{0}^{I}(\mathrm{d}x,\mathrm{d}a)$
@@
$\quad-\int_{0}^{t}\hskip-5.69046pt\int_{\mathbf{\mathbb{X}}}\psi(x,t-s)f_{x}(t-s)\overline{\mathfrak{F}}(s,x)\bar{\mu}^{S}_{s}(\mathrm{d}x)\mathrm{d}s$
$\quad+\int_{0}^{t}\hskip-5.69046pt\int_{\mathbf{\mathbb{X}}}\partial_{a}\psi(x,t-s)F^{c}_{x}(t-s)\overline{\mathfrak{F}}(s,x)\bar{\mu}^{S}_{s}(\mathrm{d}x)\mathrm{d}s\,.$
@@
$\langle\bar{\mu}^{I}_{t},\partial_{a}\psi\rangle$ $=\int_{0}^{\infty}\hskip-8.5359pt\int_{\mathbf{\mathbb{X}}}\partial_{a}\psi(x,a+t)\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\bar{\mu}_{0}^{I}(\mathrm{d}x,\mathrm{d}a)$
$\quad+\int_{0}^{t}\hskip-5.69046pt\int_{\mathbf{\mathbb{X}}}\partial_{a}\psi(x,t-s)F^{c}_{x}(t-s)\overline{\mathfrak{F}}(s,x)\bar{\mu}^{S}_{s}(\mathrm{d}x)\mathrm{d}s\,$,
@@
$\langle\bar{\mu}^{I}_{t},h\psi\rangle$ $=\int_{0}^{\infty}\hskip-8.5359pt\int_{\mathbf{\mathbb{X}}}\psi(x,a+t)\frac{f_{x}(a+t)}{F^{c}_{x}(a)}\bar{\mu}_{0}^{I}(\mathrm{d}x,\mathrm{d}a)$
$\quad+\int_{0}^{t}\hskip-5.69046pt\int_{\mathbf{\mathbb{X}}}\psi(x,t-s)f_{x}(t-s)\overline{\mathfrak{F}}(s,x)\bar{\mu}^{S}_{s}(\mathrm{d}x)\mathrm{d}s\,.$
@@
$\langle\bar{\mu}^{I}_{t},\psi\rangle=\langle\bar{\mu}^{I}_{0},\psi\rangle+\int_{0}^{t}\langle\bar{\mu}^{I}_{s},\partial_{a}\psi-h\psi\rangle ds+\int_{0}^{t}\int_{\mathbf{\mathbb{X}}}\psi(x,0)\overline{\mathfrak{F}}(s,x)\bar{\mu}^{S}_{s}(\mathrm{d}x)\mathrm{d}s\,.$ (4.12)